# Semiparametric conformal prediction

### v1

@@ 10.2 Multivariate quantile
For $d>1$, the parametric submodel is $F_{t}(s)=(1-t)F(s)+t\mathbbm{1}[{\tilde{s}}_{1}\leq s_{1},\dotsc,{\tilde{s}}_{d}\leq s_{d}]$, with $s=[s_{1},\dotsc,s_{d}]$. We proceed similarly as before, but directly in terms of the CDF to avoid dealing with multiple integrals. To simplify notation, let us first define $\Psi_{1-\alpha}(P_{t})\coloneqq g(t)$ and $\Psi_{1-\alpha}(P)\coloneqq g$.
@@
$(1-t)F(g(t))+t\mathbbm{1}[\underbrace{{\tilde{s}}_{1}\leq g(t)_{1},\dotsc,{\tilde{s}}_{d}\leq g(t)_{d}}_{\tilde{s}\preccurlyeq g(t)}]=1-\alpha.$
@@
$-F(g)+\frac{\partial F}{\partial s_{1}}\bigg{|}_{s=g}\frac{dg_{1}}{dt}+\cdots+\frac{\partial F}{\partial s_{d}}\bigg{|}_{s=g}\frac{dg_{d}}{dt}+\mathbbm{1}[\tilde{s}\preccurlyeq g]=0$
$\implies-(1-\alpha)+\frac{\partial F}{\partial s_{1}}\bigg{|}_{s=g}\frac{dg_{1}}{dt}+\cdots+\frac{\partial F}{\partial s_{d}}\bigg{|}_{s=g}\frac{dg_{d}}{dt}+\mathbbm{1}[\tilde{s}\preccurlyeq g]=0$
$\implies\frac{\partial F}{\partial s_{1}}\bigg{|}_{s=g}\frac{dg_{1}}{dt}+\cdots+\frac{\partial F}{\partial s_{d}}\bigg{|}_{s=g}\frac{dg_{d}}{dt}=(1-\alpha)-\mathbbm{1}[\tilde{s}\preccurlyeq g].$ (33)
@@
$\nabla F(g)\cdot(\lambda\nabla F(g))=(1-\alpha)-\mathbbm{1}[\tilde{s}\preccurlyeq g]$
$\implies\lambda|\nabla F(g)|^{2}=(1-\alpha)-\mathbbm{1}[\tilde{s}\preccurlyeq g]$
$\implies\lambda=\frac{(1-\alpha)-\mathbbm{1}[\tilde{s}\preccurlyeq g]}{|\nabla F(g)|^{2}},$ (35)
@@
$\psi_{P}(\tilde{s})=\frac{(1-\alpha)-\mathbbm{1}[\tilde{s}\preccurlyeq g]}{|\nabla F(g)|^{2}}\nabla F(g).\qed$
@@
$\psi_{P}(\tilde{s})=\frac{(1-\alpha)-\mathbbm{1}[\tilde{s}\leq g]}{f(g)^{2}}f(g)=\frac{(1-\alpha)-\mathbbm{1}[\tilde{s}\leq g]}{f(g)}=\frac{(1-\alpha)-(1-\mathbbm{1}[g\leq\tilde{s}])}{f(g)}=\frac{\mathbbm{1}[g\leq\tilde{s}]-\alpha}{f(g)},$
@@ 11 PATHWISE DIFFERENTIABLE ESTIMANDS
Term $(\rm b)=0$ as the influence function satisfies $P[\psi_{P}]=0$, and term $(\rm a)$ asymptotically converges to a normal around zero by the central limit theorem. As discussed in [25], the term $(\rm c)$ and $(\rm e)$ can be shown to converge to zero under certain conditions as well.
@@
In turn, a sufficient condition for $P[\left(\psi_{\hat{P}}-\psi_{P}\right)^{2}|\hat{P}]=o_{p}(1)$ is that $\hat{P}$ converges to $P$ in probability. Withholding an independent subset of the data for fitting $\hat{P}$ thus helps shrink term (c). See [44] and [70] for analyses of K-fold cross validation.
@@ 13.1 Synthetic example
The penicillin production simulator [58] is a numerical test function that takes in seven input parameters that control penicillin production—culture volume, biomass, concentration, temperature, glucose concentration, substrate feed rate, substrate feed concentration, and proton concentration—and outputs three real-valued target labels: the penicillin yield, production time, and the amount of carbon dioxide byproduct emitted. As shown in Figure 8, the target quantities are highly correlated, which lead to correlated residuals in Figure 9. The residuals also have upper tails as they most likely contain the 10% experimental errors.
@@ 14 MISSING-AT-RANDOM DATA
As $d$ grows, in many practical settings, there may not be many datapoints with all $d$ target variables observed. Using the first two targets ($d=2$) of the penicillin production data, we simulate a missing-at-random (MAR) scenario [71] in which we only observe one target if the observations for another target exceeds a certain threshold. Specifically, we generate the data $\{(X^{(i)},Y^{(i)})\}_{i=1}^{420}$ and simulate the missingness variable $\{Z^{(i)}_{1}\}_{i=1}^{420}$ with $Z_{1}=1$ if $Y_{1}$ is observed and 0 otherwise. If $Y_{1}$ is missing, then the score $S_{1}$ also cannot be evaluated. Variable $Y_{2}$, on the other hand, is always observed. We define the missingness as follows:
$Z^{(i)}_{1}=\mathbbm{1}[Y^{(i)}_{2}\leq 50],$
such that approximately 40% of $Y_{1}$ is missing. We split the 420 datapoints into 300 training and 120 calibration instances. The underlying model is a univariate Lasso regressor trained independently on only the observed labels for each target [57]. Additionally, we also simulate 500 fully observed test instances.
Only using the calibration instances for which both target labels are observed results in a biased quantile estimate, as illustrated in Figure 11. To include the calibration instances with $Y_{1}$ missing, we fit a Gaussian copula with missingness imputation [72] using the EM algorithm. As Table 4 shows, all the methods without imputation overcover due to the bias introduced by missingness, whereas our methods with the imputation copula achieve the desired coverage of 0.9 and yield the most efficient prediction sets.

### v2

@@ 10.2 Multivariate quantile
For $d>1$, the parametric submodel is $F_{t}(s)=(1-t)F(s)+t\mathds{1}[{\tilde{s}}_{1}\leq s_{1},\dotsc,{\tilde{s}}_{d}\leq s_{d}]$, with $s=[s_{1},\dotsc,s_{d}]$. We proceed similarly as before, but directly in terms of the CDF to avoid dealing with multiple integrals. To simplify notation, let us first define $\Psi_{1-\alpha}(P_{t})\coloneqq g(t)$ and $\Psi_{1-\alpha}(P)\coloneqq g$.
@@
$(1-t)F(g(t))+t\mathds{1}[\underbrace{{\tilde{s}}_{1}\leq g(t)_{1},\dotsc,{\tilde{s}}_{d}\leq g(t)_{d}}_{\tilde{s}\preccurlyeq g(t)}]=1-\alpha.$
@@
$-F(g)+\frac{\partial F}{\partial s_{1}}\bigg{|}_{s=g}\frac{dg_{1}}{dt}+\cdots+\frac{\partial F}{\partial s_{d}}\bigg{|}_{s=g}\frac{dg_{d}}{dt}+\mathds{1}[\tilde{s}\preccurlyeq g]=0$
$\implies-(1-\alpha)+\frac{\partial F}{\partial s_{1}}\bigg{|}_{s=g}\frac{dg_{1}}{dt}+\cdots+\frac{\partial F}{\partial s_{d}}\bigg{|}_{s=g}\frac{dg_{d}}{dt}+\mathds{1}[\tilde{s}\preccurlyeq g]=0$
$\implies\frac{\partial F}{\partial s_{1}}\bigg{|}_{s=g}\frac{dg_{1}}{dt}+\cdots+\frac{\partial F}{\partial s_{d}}\bigg{|}_{s=g}\frac{dg_{d}}{dt}=(1-\alpha)-\mathds{1}[\tilde{s}\preccurlyeq g].$ (33)
@@
$\nabla F(g)\cdot(\lambda\nabla F(g))=(1-\alpha)-\mathds{1}[\tilde{s}\preccurlyeq g]$
$\implies\lambda|\nabla F(g)|^{2}=(1-\alpha)-\mathds{1}[\tilde{s}\preccurlyeq g]$
$\implies\lambda=\frac{(1-\alpha)-\mathds{1}[\tilde{s}\preccurlyeq g]}{|\nabla F(g)|^{2}},$ (35)
@@
$\psi_{P}(\tilde{s})=\frac{(1-\alpha)-\mathds{1}[\tilde{s}\preccurlyeq g]}{|\nabla F(g)|^{2}}\nabla F(g).\qed$
@@
$\psi_{P}(\tilde{s})=\frac{(1-\alpha)-\mathds{1}[\tilde{s}\leq g]}{f(g)^{2}}f(g)=\frac{(1-\alpha)-\mathds{1}[\tilde{s}\leq g]}{f(g)}=\frac{(1-\alpha)-(1-\mathds{1}[g\leq\tilde{s}])}{f(g)}=\frac{\mathds{1}[g\leq\tilde{s}]-\alpha}{f(g)},$
@@ 11 PATHWISE DIFFERENTIABLE ESTIMANDS
Term $(\rm b)=0$ as the influence function satisfies $P[\psi_{P}]=0$, and term $(\rm a)$ asymptotically converges to a normal around zero by the central limit theorem. As discussed in (Hines et al.,, 2022), the term $(\rm c)$ and $(\rm e)$ can be shown to converge to zero under certain conditions as well.
@@
In turn, a sufficient condition for $P[\left(\psi_{\hat{P}}-\psi_{P}\right)^{2}|\hat{P}]=o_{p}(1)$ is that $\hat{P}$ converges to $P$ in probability. Withholding an independent subset of the data for fitting $\hat{P}$ thus helps shrink term (c). See (Van der Laan et al.,, 2011) and (Chernozhukov et al.,, 2018) for analyses of K-fold cross validation. Empirically, we find that the effect of term $(\rm c)$ is small, and choose to use as much of the data for one-time copula fitting as possible rather than use K-fold splits (which would increase the computational overhead of our method) or holding out a split (which would hurt the copula fit).
12 SPLIT ALGORITHM
We present Algorithm 2, a variant of Algorithm 1 that uses an extra split of the calibration set. Differences from Algorithm 1 are indicated in orange. One split of the calibration set is used to compute the marginal ECDFs and the other to fit the copula. This variant guarantees finite-sample distribution-free validity by Theorem 5.
Algorithm 2 Semiparametric conformal prediction, split version
3:Split $\mathcal{I}$ into $\mathcal{I}_{\rm train}$, $\mathcal{I}_{\rm cal,1}$, and $\mathcal{I}_{\rm cal,2}$
4:Train the underlying algorithm $\hat{f}$ on $\mathcal{I}_{\rm train}$
5:for $i\in\mathcal{I}_{\rm cal}$ do
6:     for $j\in[d]$ do
7:         $S^{(i)}_{j}\leftarrow V(X^{(i)}_{j},Y^{(i)}_{j},{\hat{f}})$
8:     end for
9:end for
10:for  $i\in\mathcal{I}_{\rm cal,1}$  do
11:     for $j\in[d]$ do
12:         ${\hat{F}}_{j}(s)\leftarrow$ Equation 12 $\triangleright$ Marginal ECDF
14:     end for
15:end for
16: Fit the copula ${\hat{C}}$ on $\{U^{(i)}\}_{i\in\mathcal{I}_{\rm cal,2}}$
17:$U^{*}\leftarrow$ Equation 16
18:$Q_{1-\alpha}\leftarrow Q_{1-\alpha}(U^{*})$ from Equation 17
19:if correction is True then
20:     $U^{*}\leftarrow U_{\rm 1\shorthyphen step}$ from Equation 19
21:     $Q_{1-\alpha}\leftarrow Q_{1-\alpha}(U^{*})$ from Equation 17
22:end if
23:for $i\in\mathcal{I}_{\rm test}$ do
24:     $\Gamma_{1-\alpha}(X^{(i)})\leftarrow$ Equation 6
25:end for
13 THEORETICAL RESULTS
Theorem 2.
Proof.
Define the $(1-\alpha)$-quantile of $F^{\ast}$ as
$Q_{1-\alpha}=\inf_{s}\{||s||_{1}:F^{\ast}(s)=1-\alpha\},$
and let $\hat{Q}_{1-\alpha}$ be the plug-in estimator obtained from $\hat{F}$. Since the quantile functional $\Psi:F\mapsto Q_{1-\alpha}$ is Hadamard differentiable under the stated regularity conditions, we have
$\sqrt{n}\Bigl{(}\hat{Q}_{1-\alpha}-Q_{1-\alpha}\Bigr{)}=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\psi(S^{(i)})+o_{p}(1),$
where $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the EIF for the quantile functional. The one-step corrected estimator is defined by
$\hat{Q}_{1-\alpha}^{\text{1-step}}=\hat{Q}_{1-\alpha}+\frac{1}{n}\sum_{i=1}^{n}\psi(S^{(i)}).$
It then follows that
$\sqrt{n}\Bigl{(}\hat{Q}_{1-\alpha}^{\text{1-step}}-Q_{1-\alpha}\Bigr{)}=o_{p}(1),$
so that $\hat{Q}_{1-\alpha}^{\text{1-step}}$ is a $\sqrt{n}$-consistent estimator of $Q_{1-\alpha}$.
Now consider a test point $(X^{\ast},Y^{\ast})$ with associated nonconformity score $S^{\ast}=V(X^{\ast},Y^{\ast},\hat{f}).$ Since the prediction set is defined by
$\Gamma_{1-\alpha}(X^{\ast})=\{Y\in\mathbb{R}^{d}:V(X^{\ast},Y,\hat{f})\preceq\hat{Q}_{1-\alpha}^{\text{1-step}}\},$
we have
$\mathbb{P}\bigl{(}Y^{\ast}\in\Gamma_{1-\alpha}(X^{\ast})\bigr{)}=\mathbb{P}\bigl{(}S^{\ast}\preceq\hat{Q}_{1-\alpha}^{\text{1-step}}\bigr{)}.$
By the definition of $\hat{F}$ and using the uniform consistency of $\hat{F}$ together with the continuity of $F^{\ast}$, it holds that
[…]
Thus, as $n\to\infty$,
$\mathbb{P}\bigl{(}S^{\ast}\preceq\hat{Q}_{1-\alpha}^{\text{1-step}}\bigr{)}=\hat{F}\bigl{(}\hat{Q}_{1-\alpha}^{\text{1-step}}\bigr{)}\to 1-\alpha.$
Theorem 3.
Proof.
We assumed that
$\sup_{S}\Bigl{|}F^{*}(S)-\hat{F}(S)\Bigr{|}\leq\epsilon.$
Let ${\hat{Q}}_{1-\alpha}$ be defined as the $(1-\alpha)$-quantile of $\hat{F}$, i.e.,
$\hat{F}({\hat{Q}}_{1-\alpha})=1-\alpha.$
Recall that the prediction set using the plug-in quantile is defined as
$\Gamma_{1-\alpha}(X^{*})=\{Y\in\mathbb{R}^{d}:V(X^{*},Y,\hat{f})\preccurlyeq{\hat{Q}}_{1-\alpha}\}.$
Note that the mapping from the distribution $F$ to the quantile is Lipschitz continuous under the TV distance. More precisely, by the data processing inequality, which states that for any measurable function $g$,
$d_{\mathrm{TV}}\bigl{(}P\circ g^{-1},Q\circ g^{-1}\bigr{)}\leq d_{\mathrm{TV}}(P,Q),$
we have for the quantile
$A(F^{\prime})=\inf_{s}\{||s||_{1}:F^{\prime}=1-\alpha\},$
that
$\Bigl{|}A(F)-A(\hat{F})\Bigr{|}\leq\sup_{S}\Bigl{|}F^{*}(S)-\hat{F}(S)\Bigr{|}\leq\epsilon.$
Since $\hat{F}({\hat{Q}}_{1-\alpha})=1-\alpha$, it follows that
$F^{*}({\hat{Q}}_{1-\alpha})\geq 1-\alpha-\epsilon.$
Now the coverage is
$\mathbb{P}\Bigl{[}Y^{*}\in\Gamma_{1-\alpha}(X^{*})\Bigr{]}=\mathbb{P}\Bigl{[}V(X^{*},Y^{*},\hat{f})\leq{\hat{Q}}_{1-\alpha}\Bigr{]}=F^{*}({\hat{Q}}_{1-\alpha})\geq 1-\alpha-\epsilon.$
as desired. If instead we use the one-step corrected quantile to form the prediction set $\Gamma_{1-\alpha}^{\text{1-step}}$, then by design the correction term does not introduce additional error beyond the estimation error of $\hat{F}$. Hence, the same argument applies, and we again obtain
$\mathbb{P}\Bigl{[}Y^{*}\in\Gamma_{1-\alpha}^{\text{1-step}}(X^{*})\Bigr{]}\geq 1-\alpha-\epsilon.$
Theorem 5 (Finite-sample validity of Algorithm 2).
Assume that the full dataset (including calibration data and the test point) is exchangeable, and that the calibration set is split into two disjoint subsets,
$I_{\mathrm{cal}}=I_{\mathrm{cal},1}\cup I_{\mathrm{cal},2},$
where the set $I_{\mathrm{cal},1}$ is used to construct the marginal empirical CDFs and to fit the copula, and the set $I_{\mathrm{cal},2}$ (with $n^{\prime}\equiv|I_{\mathrm{cal},2}|$) is used to construct the empirical copula, from which the $\lceil(1-\alpha)(n^{\prime}+1)\rceil/(n^{\prime}+1)$-quantile is extracted. Then the prediction set
$\Gamma_{1-\alpha}(X^{*})=\{y\in\mathbb{R}^{d}:V(X^{*},y,\hat{f})\preceq Q_{1-\alpha}\},$
constructed with the quantile $Q_{1-\alpha}$ from $I_{\mathrm{cal},2}$ satisfies
$\mathbb{P}\Bigl{(}Y^{*}\in\Gamma_{1-\alpha}(X^{*})\Bigr{)}\geq 1-\alpha.$
Proof.
Since the data are exchangeable, the procedure of splitting the calibration set into $I_{\mathrm{cal},1}$ and $I_{\mathrm{cal},2}$ guarantees that the scores computed on $I_{\mathrm{cal},2}$ are independent of the marginal and copula estimates that were obtained from $I_{\mathrm{cal},1}$. Denote by
$S_{1},S_{2},\ldots,S_{n^{\prime}}$
the scores computed on the samples in $I_{\mathrm{cal},2}$, and let $S^{*}$ be the score computed on the test point $(X^{*},Y^{*})$. By construction, the scores $S_{1},\ldots,S_{n^{\prime}}$ and $S^{*}$ are exchangeable.
Let $Q_{1-\alpha}$ be the $\lceil(1-\alpha)(n^{\prime}+1)\rceil$-th order statistic of the $n^{\prime}$ calibration scores. By the standard conformal prediction argument (see, e.g., Vovk et al., (2005); Lei and Wasserman, (2014)), the rank of $S^{*}$ among the $n^{\prime}+1$ scores is uniformly distributed. We thus have
$\mathbb{P}\Bigl{(}S^{*}\preceq Q_{1-\alpha}\Bigr{)}\geq\frac{\lceil(n^{\prime}+1)(1-\alpha)\rceil}{n^{\prime}+1}\geq 1-\alpha.$
Since the prediction set is defined as
$\Gamma_{1-\alpha}(X^{*})=\{y\in\mathbb{R}^{d}:V(X^{*},y,\hat{f})\preceq Q_{1-\alpha}\},$
it follows that
$\mathbb{P}\Bigl{(}Y^{*}\in\Gamma_{1-\alpha}(X^{*})\Bigr{)}=\mathbb{P}\Bigl{(}S^{*}\preceq Q_{1-\alpha}\Bigr{)}\geq 1-\alpha.$
@@ 14 CALIBRATION SCHEMES
Table 4 includes a full set of experiments for the penicillin data ($d=3$) with a bigger calibration set ($n=200$), to accommodate methods requiring an extra split of the calibration set. Too few calibration points for the second split $\mathcal{I}_{{\rm cal},2}$ will result in infinitely large prediction sets. These methods are labeled “split.”
For scalar score methods using an extra split, the first split $\mathcal{I}_{{\rm cal},1}$ is used to compute the mean and standard deviation of the per-target scores with which to standardize the scores of $\mathcal{I}_{{\rm cal},2}$, for usual ranking. This method can be viewed as using a “fitted target aggregation scheme,” but still carries the fundamental limitation of scalar scores, of not accounting for the joint correlation structure. Moreover, using the calibration split only to standardize the scores directly may be a waste when the labels can be standardized everywhere, as was done already, at no cost. We observe that the split scalar scores still overcover, as they do not perform joint ranking.
Table 4: Metrics on the penicillin data ($d=3$, $n=200$). We report mean $\pm$ standard error across ten seeds. Target coverage is 0.9. For efficiency, lower is better.
Method Coverage Efficiency $\downarrow$
Independent $0.95\pm 0.00$ $4.0\pm 0.1$
Scalar score ($L_{2}$) $0.92\pm 0.00$ $4.3\pm 0.1$
Scalar score ($L_{1}$) $0.92\pm 0.00$ $4.6\pm 0.1$
Scalar score ($L_{\infty}$) $0.92\pm 0.00$ $4.0\pm 0.1$
Empirical copula $0.91\pm 0.01$ $3.6\pm 0.1$
Scalar score ($L_{2}$), split $0.93\pm 0.01$ $4.4\pm 0.1$
Scalar score ($L_{1}$), split $0.93\pm 0.01$ $4.8\pm 0.2$
Scalar score ($L_{\infty}$), split $0.93\pm 0.01$ $4.0\pm 0.1$
Plug-in (ours) $0.92\pm 0.00$ $3.6\pm 0.2$
Corrected (ours) $0.90\pm 0.00$ $3.6\pm 0.1$
Plug-in, split (ours) $0.92\pm 0.03$ $3.8\pm 0.2$
Corrected, split (ours) $0.92\pm 0.03$ $3.8\pm 0.2$
@@ 15.1 Synthetic example
The penicillin production simulator (Liang and Lai,, 2021) is a numerical test function that takes in seven input parameters that control penicillin production—culture volume, biomass, concentration, temperature, glucose concentration, substrate feed rate, substrate feed concentration, and proton concentration—and outputs three real-valued target labels: the penicillin yield, production time, and the amount of carbon dioxide byproduct emitted. As shown in Figure 8, the target quantities are highly correlated, which lead to correlated residuals in Figure 9. The residuals also have upper tails as they most likely contain the 10% experimental errors.
@@
Figure 10 plots the mean and standard deviation in coverage for all the methods across ten random seeds as the number of calibration points $n$ increases. Our corrected method consistently yields valid coverage at the 0.91-0.92 level for a target coverage of 0.9. At small $n$, our ”plug-in” method has mean coverage of 0.91 with high std (0.1) across ten random seeds, but one-step correction appropriately adjusts the coverage upward. For all $n$, the variances of coverage for our methods are similar to the those of “independent” and “scalar score” baselines for all and shrink to 0.04 at $n\sim 200$.
Figure 10: One-step correction improves the coverage of the plug-in estimate at all calibration-set sizes $n$ for the penicillin ($d=3$) dataset. Error bars are stddev across ten random seeds.
@@ 16 MISSING-AT-RANDOM DATA
As $d$ grows, in many practical settings, there may not be many datapoints with all $d$ target variables observed. Using the first two targets ($d=2$) of the penicillin production data, we simulate a missing-at-random (MAR) scenario (Rubin,, 1976) in which we only observe one target if the observations for another target exceeds a certain threshold. Specifically, we generate the data $\{(X^{(i)},Y^{(i)})\}_{i=1}^{420}$ and simulate the missingness variable $\{Z^{(i)}_{1}\}_{i=1}^{420}$ with $Z_{1}=1$ if $Y_{1}$ is observed and 0 otherwise. If $Y_{1}$ is missing, then the score $S_{1}$ also cannot be evaluated. Variable $Y_{2}$, on the other hand, is always observed. We define the missingness as follows:
$Z^{(i)}_{1}=\mathds{1}[Y^{(i)}_{2}\leq 50],$
such that approximately 40% of $Y_{1}$ is missing. We split the 420 datapoints into 300 training and 120 calibration instances. The underlying model is a univariate Lasso regressor trained independently on only the observed labels for each target (Tibshirani,, 1996). Additionally, we also simulate 500 fully observed test instances.
Only using the calibration instances for which both target labels are observed results in a biased quantile estimate, as illustrated in Figure 12. To include the calibration instances with $Y_{1}$ missing, we fit a Gaussian copula with missingness imputation Feldman and Kowal, (2024) using the EM algorithm. As Table 5 shows, all the methods without imputation overcover due to the bias introduced by missingness, whereas our methods with the imputation copula achieve the desired coverage of 0.9 and yield the most efficient prediction sets.